Raindrops keep fallin' on your head. Assuming the drops are 2mm in diameter and your head is struck by 12 of them per second, what is the average force on your head in N ? Note that this problem requires you think about momentum and several other concepts from the class. Hint: the coefficient of air drag is
1
Expert's answer
2020-09-16T10:09:28-0400

"1N=9m\/s" Thus "12m\/s =1.3N= 12.74kg"


"V=10^{-3}m m=12.74kg"


we know that: "V^2-U^2=2as"


Thus "0-(12)=2a\times 10^{-3}"

"a=-144\times 10^3m/s^2"

But "F=ma"


"12.74\times (-144\times10^3)= 18.3N"


Average force= 18.3N